How much water must be added to 150 liters of denatured alcohol 85% pure to dilute it into 75% alcohol content?
Answer by CharlesG2(834)

How much water must be added to 150 liters of denatured alcohol 85% pure to dilute it into 75% alcohol content
100% - 85% = 15%, 100% - 75% = 25 %
% in A * amount A + % in B * amount B = % in C * amount C
100% * x + 15% * 150 = 25% * (x + 150)
x + 0.15 * 150 = 0.25x + 0.25 * 150
x + 22.5 = 0.25x + 37.5
0.75x = 15
x = 20 liters